(a) Let P(x) be the statement x
2 ≥ x.
(i) What are truth values of the propositions P(1) , P(-1) , P(0) and P( 1
2
)?
(ii) What is the truth value of the proposition ∀xP(x), where the domain consists of all
real numbers?
(iii) What is the truth value of the proposition ∀xP(x), where the domain consists of all
integers?
(b) Let Q(x,y,z) be the statement x
2 + y
2 = z
2
. What are the truth values of
(i) ∀(x,y,z) Q(x,y,z)
(ii) ∃(x,y,z) Q(x,y,z) Where the domain consists of all real numbers.
1
Expert's answer
2020-06-11T19:37:20-0400

a) "P(x):x^2\geq x"

i)

"P(1): 1=1" - true

"P(-1): 1>-1" - true

"P(0): 0=0" - true

"P(1\/2): 1\/4<1\/2" - false


ii)

Answer: false (example: "P(1\/2): 1\/4<1\/2" )


iii)

Answer: true


b) "Q(x,y,z):x^2+y^2=z^2"

i)

Answer: false


ii)

Answer: true (example: "3^2+4^2=5^2" )